# Cosmological constant as an integration constant

### v2

@@ Abstract
The discrepancy between the observed value of the cosmological constant (CC) and its expected value from quantum field theoretical considerations motivates the search for a theory in which the CC is decoupled from the vacuum energy. In this article, we consider theories in which the CC is regarded as an integration constant. These theories include trace-free Einstein gravity, theories constructed from the Codazzi equation (which includes Cotton gravity and a gauge-gravity inspired theory), and conformal Killing gravity. We remark on a recent debate regarding the viability of Cotton gravity and find that while the Codazzi equation is indeed underdetermined, the solutions of the Codazzi equation trivialize to $\lambda g_{ab}$ on generic backgrounds, and that in principle, one can close the system with the divergence-free condition and an appropriate choice of initial data. We also propose a full variational principle (full in the sense that variations in all variables are considered) for each of the aforementioned theories that can incorporate the matter sector; in this manner, we can obtain the trace-free Einstein equations without the unimodular constraint.
@@
Such a reformulation should somehow decouple the CC from microscopic physics. One strategy is to seek a formulation of Einstein gravity in which the CC arises as a constant of integration. At the level of the field equations, there are several proposals in the literature. A proposal originating from Einstein [5] is that of the trace-free Einstein equations obtained in unimodular gravity [6, 4]. Another proposal by Cook and Chen [7, 8], inspired by a set of third-order equations developed by Kilmister, Newman [9, *Kilmister:1961], and later by Yang [11] is to construct a gravity theory from the Bianchi identities and derivatives of the energy-momentum tensor. More recently, Harada proposed two theories of gravity, termed Cotton gravity [12] and the so-called conformal Killing gravity [13, 14]. Cotton gravity is constructed from the Cotton tensor, and it has been shown that one can rewrite the field equations of Cotton gravity as a Codazzi equation $\nabla_{[a}C_{b]c}=0$ [15, 16]; in this regard, Cotton gravity is in the same class as the earlier theory of Cook and Chen, as we will show that the field equations for the latter can also be reformulated as a Codazzi equation. conformal Killing gravity, on the other hand, has a different structure altogether, as the field equations have the form of the defining equation for a gradient conformal Killing tensor.
@@
Cotton gravity has drawn some criticism [17] and an extended debate[18, 19, 20] over its viability as a physical theory, as it was discovered that the equations of Cotton gravity are underdetermined under a symmetry reduction. [^1] It is perhaps worth mentioning that the equations of Kilmister, Newman, and Yang, that inspired the approach of Cook and Chen, also invited a flurry of debate [21, 22, 23] shortly after the publication of [11], and some discussion regarding the physicality of its solutions [24, 25]. It is been argued in [18, 20] that while the equations of Cotton gravity are indeed underdetermined in cases of high symmetry (for instance, cosmological or spherically symmetric spacetimes), this is not necessarily the case in generic spacetimes. Similar concerns apply to the theory of Cook and Chen, as they share the same structure. Another concern, that applies to the theories of Cook and Chen, Cotton gravity, and conformal Killing gravity, is the fact that they are only defined at the level of the field equations; all three theories suffer from the lack of a complete variational principle, as the variational principles proposed so far require holding some of the variables (such as the spacetime metric $g_{ab}$) fixed. One difficulty is that the energy-momentum tensor is obtained from variations of the metric tensor in the matter sector, but derivatives of the energy-momentum tensor appear in the field equations.
@@
Later, we will consider a variational principle that does not impose a unimodular constraint on the metric.
@@
The tensor $X_{\mu\nu}$ is assumed to be determined by boundary and initial data; if the initial/boundary data yields $X_{\mu\nu}=\lambda g_{\mu\nu}$, then the CC $\lambda$ may be interpreted as an integration constant (we will discuss later how such a solution arises from initial data).
@@
where $u_{a}$ is a one-form; $K_{ab}$ is a symmetric Killing tensor when $u_{a}=0$. However, as pointed out in [27] a simple redefinition $\bar{K}_{ab}:=6K_{ab}-g_{ab}K$ reveals that any nontrivial solution of Eq. (16) requires the existence of a Killing tensor given by $\bar{K}_{ab}$. It follows that the solutions of Eq. (16) either trivialize to $K_{ab}=-\lambda g_{ab}$, or admit a Killing tensor. For this reason, it is perhaps more appropriate to refer to this theory as “Killing tensor gravity”, but for the sake of clarity, we retain the more common terminology for this theory. Moreover, we note that perturbations of Killing tensor solutions satisfying $K_{ab}\neq-\lambda g_{ab}$ must also admit a Killing tensor; since our universe (being inhomogeneous and anisotropic on small scales) does not have the required symmetries to admit a global Killing tensor, we regard such solutions as unphysical.
@@
Note that the right-hand side contains at most first derivatives of $C_{ab}$ (and also that the divergence-free condition removes the first term). One can then regard Eq. (24) as a wave equation for $Q_{cde}$, with $C_{ab}$ being an independent variable satisfying Eq. (20) and the constraint (24). Now consider initial data consistent with Eq. (18), so that $Q_{cde}=0$ (but not its derivatives). However, for nonvanishing curvature, the right-hand side will generically generate a source unless both $C_{ab}$ and all its first derivatives vanish. This indicates that if $C_{ab}$ is nontrivial, then even for initial data satisfying Eq. (24) on a Cauchy surface, the time evolution of $C_{ab}$ under Eq. (20) will in general generate violations of Eq. (24). This can be avoided if $C_{ab}=\lambda g_{ab}$, which satisfies Eq. (20), and yields a vanishing right-hand side for Eq. (24) if one also sets $Q_{cde}=0$ at the initial surface. There may also be special cases (highly symmetric spacetimes, for instance) in which Eq. (18) may admit nontrivial solutions, but one cannot expect the existence of solutions to Eq. (18) for nontrivial initial data. It should be noted that Eq. (24) and the conclusions outlined in the previous paragraph are independent of the constraint $\nabla_{a}C^{ab}=0$.
@@
Variational principle — Finding a satisfactory variational principle for the aforementioned field equations is a challenge. The variational principles presented in the literature for the trace-free Einstein equations [4] and theories admitting a Codazzi equation formulation (Bianchi gravity [7, 8] and Cotton gravity [12]) typically require holding certain quantities fixed; for the trace-free Einstein equations, the metric determinant is held fixed (corresponding to an additional constraint), and for the Codazzi equation, the metric is held fixed (as variations with respect to the metric will generally introduce additional unwanted constraints). We therefore seek a general variational principle that yields the desired field equations (including contributions from the matter sector) without unwanted constraints under a full variation—that is, a variation with respect to all variables appearing in the action.
Consider a simple action of the form for the case of the Codazzii equation:
@@
which reduces to the Codazzi equation when $Y^{abc}=0$. The variation of the action $S_{C}$ with respect to $g^{ab}$ and $Z_{ab}$ is rather nontrivial, but will ultimately consist of terms proportional to products of $Y^{abc}$ and its derivatives; the variations with respect to $g^{ab}$ and $Z_{ab}$ vanish when $Y^{abc}=0$. If one wishes to independently enforce the divergence-free constraint, one may add an action of the form:
@@
Discussion — In this article, we discussed theories in which the CC arises as an integration constant, including trace-free Einstein gravity, a class of theories that can be formulated as a Codazzi equation, which includes Cotton gravity [12] and a theory developed by Cook and Chen [7, 8], and conformal Killing gravity. We revisited a recent debate surrounding the issue of nonuniqueness of solutions to the equations of Cotton gravity [17, 18, 19, 20], particularly in high-symmetry situations of physical interest. By taking additional derivatives of the Codazzi equations, we find that solutions of the Codazzi equations must satisfy at least one of two conditions: either the solutions coincide with that of the Einstein equations (with arbitrary CC), or that certain contractions of the curvature tensor, the Codazzi tensor, and their derivatives must vanish (in particular, all terms on the right-hand side of Eq. (24) must vanish). The latter can be satisfied in situations of high symmetry, even if the solutions do not satisfy the Einstein equations, but if we are to regard such non-Einstein solutions as physical, the right-hand side of Eq. (24) has to vanish exactly even under perturbations. It is unlikely that such conditions can be satisfied unless the perturbations are highly restricted, and one may in this manner argue that such non-Einstein solutions are unphysical. We have also found that the Codazzi equations can in principle be closed with the addition of the divergence-free constraint so that such non-Einstein solutions can be excluded at the level of initial data.
@@
We also formulated a rather general variational principle that can yield the trace-free Einstein equations, the Codazzi equation, and the equations of conformal Killing gravity. An action for trace-free gravity without a unimodular constraint [28] was recently proposed, but the matter sector was not explicitly incorporated in their treatment; our action (30) provides an alternative that can readily incorporate the matter sector via the action (32). Now one might worry about the use of Lagrange multipliers and auxiliary fields in our variational principle as one is effectively introducing additional degrees of freedom—one might prefer an action that depends exclusively on dynamical degrees of freedom. Moreover, there is some question of whether auxiliary fields in gravity theories are pathological [29, 30, 31]. Regarding the latter, there may be instances in which some of the pathologies (particularly those associated with sharp density gradients) can be mitigated [32, 33, 34], and one can perhaps regard auxiliary fields as effective descriptions for underlying dynamical degrees of freedom (with large masses). Regarding the former concern, it is well-known that one can express the Lagrangian for $f(R)$ theories in the O’Hanlon [35] form $\psi R-V(\psi)$ (see also [36, 37, 38, 39, 40] and references therein), where $\psi$ is an auxiliary field, and one can easily recover the $f(R)$ Lagrangian from the O’Hanlon Lagrangian by algebraic elimination. Of course, a naive algebraic elimination would trivialize the actions considered here, but the point is that auxiliary fields can provide a starting point for constructing an action exclusively from dynamical degrees of freedom. To clarify, we are not suggesting that the variational principle we have proposed is fundamental—indeed, one can construct a variational principle for virtually any set of consistent field equations in this manner. [^4] Presumably one could attempt to construct a variational principle for inconsistent field equations, but this will result in an action that does not admit extrema. Instead, we argue that it can serve as a first step towards the construction of a more fundamental action for these theories.
@@
One might be concerned about ghosts, as the resulting equations contain up to three derivatives of the metric. One might also recognize that the duplicate matter degrees of freedom $\chi$ in Eq. (32) are necessarily ghosts. At the classical level, there is no issue in the matter sector, as the matter field equations ultimately reduce to those identical to that obtained in the standard variational principle—there is no runaway behavior (see also [41]). In the gravitational sector, one can place restrictions on the initial data such that the system becomes dynamically equivalent to the Einstein equations with a CC. Regarding the Hamiltonian, we note that on shell, both the matter and gravitational actions vanish, and it is not too difficult to convince oneself that the total matter Hamiltonian must also vanish on shell. Though there is no issue at the classical level, the appearance of ghosts may be of particular concern for the quantization of the theory. Quantization is beyond the scope of the present article (see the discussion in [8] regarding ghosts in higher curvature theories), but we leave the reader with a couple of remarks. First, as argued in the preceding paragraph, the action presented in this article may require further modification at high energies, and the structure of the resulting action may differ significantly. Second, the vanishing of the Hamiltonian indicates that these theories suffer from the problem of time, in which Hamiltonian operator does not yield the time evolution of the quantum state of the field, a well-known problem in canonical approaches to quantum gravity [42, *Isham:1992ms].
@@
Acknowledgments — We thank Marco de Cesare and Naresh Dadhich for helpful comments. PC and JCF acknowledge support from the Taiwan National Science and Technology Council (NSTC) under project numbers 110-2112-M-002-031 and 112-2811-M-002-132, and also from the Leung Center for Cosmology and Particle Astrophysics (LeCosPA), National Taiwan University.

### v3

@@ Abstract
The discrepancy between the observed value of the cosmological constant (CC) and its expected value from quantum field theoretical considerations motivates the search for a theory in which the CC is decoupled from the vacuum energy. In this article, we consider theories in which the CC is regarded as an integration constant. These theories include trace-free Einstein gravity, theories constructed from the Codazzi equation (which includes Cotton gravity and a gauge-gravity inspired theory), and conformal Killing gravity. We remark on a recent debate regarding the viability of Cotton gravity and find that while the Codazzi equation is indeed underdetermined, the solutions of the Codazzi equation trivialize to $\lambda g_{ab}$ on generic backgrounds, and that in principle, one can close the system with the divergence-free condition and an appropriate choice of initial data. We also propose a full variational principle (full in the sense that variations in all variables are considered) for each of the aforementioned theories that can incorporate the matter sector; in this manner, we can obtain the trace-free Einstein equations without a unimodular constraint.
@@
Such a reformulation should somehow decouple the CC from microscopic physics. One strategy is to seek a formulation of Einstein gravity in which the CC arises as a constant of integration. At the level of the field equations, there are several proposals in the literature. A proposal originating from Einstein [5] is that of the trace-free Einstein equations, which are obtained in unimodular gravity [6, *vanderBij:1981ym, *Buchmuller:1988wx, *Buchmuller:1988yn, *Henneaux:1989zc, 4] (see [11, *Alvarez:2023utn] for recent overviews of the topic), but can be considered on their own without the constraint on the metric determinant [13, 14]. Another proposal by Cook and Chen [15, 16], inspired by a set of third-order equations developed by Kilmister, Newman [17, *Kilmister:1961], and later by Yang [19] is to construct a gravity theory from the Bianchi identities and derivatives of the energy-momentum tensor. We focus primarily on the aforementioned approaches, though we mention here other proposals in which one may obtain the unconstrained Einstein field equations with the CC arising as an integration constant, such as the thermodynamical approach of Padmanabhan [20, 21] (see also [22]), a modified metric-affine theory [23, *Tapia:1998qt, *Tapia:1998as], and contributions from a three-form action  [^1] However, we note that in this case the vacuum energy still contributes to the cosmological term. [26, 27]. More recently, Harada proposed two theories of gravity, termed Cotton gravity [28] and the so-called conformal Killing gravity [29, 30]. Cotton gravity is constructed from the Cotton tensor, and it has been shown that one can rewrite the field equations of Cotton gravity as a Codazzi equation $\nabla_{[a}C_{b]c}=0$ [31, 32]; in this regard, Cotton gravity is in the same class as the earlier theory of Cook and Chen, as we will show that the field equations for the latter can also be reformulated as a Codazzi equation. Conformal Killing gravity, on the other hand, has a different structure altogether, as the field equations have the form of the defining equation for a gradient conformal Killing tensor.
@@
Cotton gravity was recently the focus of a debate [33, 34, 35, 36] over its viability as a physical theory, as it was discovered that the symmetry-reduced equations of Cotton gravity are underdetermined. It is been argued in [34, 36] that while the equations of Cotton gravity are indeed underdetermined in cases of high symmetry (for instance, cosmological or spherically symmetric spacetimes), this is not necessarily the case in generic spacetimes. Similar concerns apply to the theory of Cook and Chen, as they share the same Codazzi structure. Another concern that applies to Codazzi type theories and conformal Killing gravity, is the fact that they are only defined at the level of the field equations; these theories suffer from the lack of a complete variational principle, as the variational principles proposed so far require holding some of the variables (such as the spacetime metric $g_{ab}$) fixed. One difficulty is that the energy-momentum tensor is obtained from variations of the metric tensor in the matter sector, but derivatives of the energy-momentum tensor appear in the field equations. The primary aim of this article is to shed light on the debate over the Codazzi equation and to propose a variational principle (including matter) for the trace-free Einstein equations, the Codazzi equation, and the conformal Killing equation without introducing additional constraints.
@@
Later, we consider a variational principle that does not impose a unimodular constraint on the metric.
@@
The tensor $X_{\mu\nu}$ is assumed to be determined by boundary and initial data; if the initial/boundary data yields $X_{\mu\nu}=\lambda g_{\mu\nu}$, then the CC $\lambda$ may be interpreted as an integration constant (we discuss later how such a solution arises from initial data).
@@
where $u_{a}$ is a one-form; $K_{ab}$ is a symmetric Killing tensor when $u_{a}=0$. However, as pointed out in [37] a simple redefinition $\bar{K}_{ab}:=6K_{ab}-g_{ab}K$ reveals that any nontrivial solution of Eq. (16) requires the existence of a Killing tensor given by $\bar{K}_{ab}$. It follows that the solutions of Eq. (16) either trivialize to $K_{ab}=-\lambda g_{ab}$, or admit a Killing tensor. For this reason, it is perhaps more appropriate to refer to this theory as “Killing tensor gravity”, but we retain here the common terminology in the literature. Moreover, we note that perturbations of Killing tensor solutions satisfying $K_{ab}\neq-\lambda g_{ab}$ must also admit a Killing tensor; since our universe (being inhomogeneous and anisotropic on small scales) does not have the required symmetries to admit a global Killing tensor, we regard such solutions as unphysical.
@@
Note that the right-hand side contains at most first derivatives of $C_{ab}$ (and also that the divergence-free condition removes the first term). One can then regard Eq. (24) as a wave equation for $Q_{cde}$, with $C_{ab}$ being an independent variable satisfying Eq. (20) and the constraint (24). Now consider initial data consistent with Eq. (18), so that $Q_{cde}=0$ (but not its derivatives). For nonvanishing curvature, the right-hand side will generically produce a source unless $C_{ab}=\lambda g_{ab}$ on a Cauchy hypersurface and $\nabla_{c}C_{ab}=0$. This indicates that if $C_{ab}\neq\lambda g_{ab}$ and $\nabla_{c}C_{ab}\neq 0$ satisfies Eq. (24) on a Cauchy surface, the time evolution of $C_{ab}$ under Eq. (20) will in general generate violations of Eq. (24). There may also be special cases (highly symmetric spacetimes, for instance) in which Eq. (18) may admit nontrivial solutions $C_{ab}\neq\lambda g_{ab}$, but one cannot expect the existence of solutions to Eq. (18) for generic initial data. It should be noted that Eq. (24) and the conclusions outlined in the previous paragraph are independent of the constraint $\nabla_{a}C^{ab}=0$.
@@
Variational principle — The variational principles presented in the literature for the trace-free Einstein equations [4] and theories admitting a Codazzi equation formulation (Bianchi gravity [15, 16] and Cotton gravity [28]) typically require holding certain quantities fixed; for the trace-free Einstein equations, the metric determinant is held fixed (corresponding to an additional constraint), and for the Codazzi equation, the metric is held fixed (as variations in the metric will generally introduce additional unwanted constraints). We therefore seek a general variational principle that yields the desired field equations (including contributions from the matter sector) without unwanted constraints under a full variation—that is, a variation with respect to all variables appearing in the action.
Consider a simple action of the form for the case of the Codazzi equation:
@@
which reduces to the Codazzi equation when $Y^{abc}=0$. The variation of the action $S_{C}$ with respect to $g^{ab}$ and $Z_{ab}$ is rather nontrivial, but ultimately consists of terms proportional to products of $Y^{abc}$ and its derivatives; variations with respect to $g^{ab}$ and $Z_{ab}$ vanish when $Y^{abc}=0$. To independently enforce the divergence-free constraint, one may add to the action the following:
@@
Discussion — In this article, we considered some theories in which the CC arises as an integration constant, including trace-free Einstein gravity, the Codazzi equation, and conformal Killing gravity. We revisited a recent debate [33, 34, 35, 36] concerning the issue of nonuniqueness of solutions to equations of Cotton gravity in symmetry-reduced settiongs. By taking additional derivatives of the Codazzi equations, we find that solutions of the Codazzi equations must satisfy at least one of two conditions: either the solutions coincide with that of the Einstein equations (with arbitrary CC), or that certain contractions of the curvature tensor, the Codazzi tensor, and their derivatives must vanish (in particular, all terms on the right-hand side of Eq. (24) must vanish). The latter can be satisfied in situations of high symmetry, even if the solutions do not satisfy the Einstein equations, but if we regard such non-Einstein solutions as physical, the right-hand side of Eq. (24) must vanish exactly even under perturbations. It is unlikely that such conditions can be satisfied unless the perturbations are highly restricted, and one may in this manner argue that such non-Einstein solutions are unphysical. We find that the Codazzi equations can in principle be closed with the addition of the divergence-free constraint so that such non-Einstein solutions can be excluded at the level of initial data.
@@
We formulated a rather general variational principle that can yield the trace-free Einstein equations, the Codazzi equation, and the equations of conformal Killing gravity. An action for trace-free gravity without a unimodular constraint [38] was recently proposed (another possible approach is described in [23, *Tapia:1998qt, *Tapia:1998as]), but the matter sector was not explicitly incorporated in their treatment; our action (30) provides an alternative that can readily incorporate the matter sector via the action (32). The appearance of Lagrange multipliers and auxiliary fields in the actions corresponds to introducing additional nondynamical degrees of freedom—one might prefer an action that depends exclusively on dynamical degrees of freedom. Moreover, there is some question of whether auxiliary fields in gravity theories are pathological [39, 40, 41]. Regarding the latter, there may be instances in which some of the pathologies (particularly those associated with sharp density gradients) can be mitigated [42, 43, 44], and one can perhaps regard auxiliary fields as effective descriptions for underlying dynamical degrees of freedom (with large masses). Regarding the former concern, it is well-known that one can express the Lagrangian for $f(R)$ theories in the O’Hanlon [45] form $\psi R-V(\psi)$ (see also [46, 47, 48, 49, 50] and references therein), where $\psi$ is an auxiliary field, and one can easily recover the $f(R)$ Lagrangian from the O’Hanlon Lagrangian by algebraic elimination. Of course, a naive algebraic elimination would trivialize the actions considered here, but the point is that auxiliary fields can provide a starting point for constructing an action exclusively from dynamical degrees of freedom. To clarify, we are not suggesting that the variational principle we have proposed is fundamental—indeed, one can construct a variational principle for virtually any set of consistent field equations in this manner. [^4] Presumably one could attempt to use the same strategy to construct a variational principle for inconsistent field equations, but this will result in an action that does not admit extrema. Instead, we argue that it can serve as a first step towards the construction of a more fundamental action for these theories.
@@
Ghosts are a concern; for instance, the Codazzi equation contains up to three derivatives of the metric. One might also recognize that the duplicate matter degrees of freedom $\chi$ in Eq. (32) are necessarily ghosts. At the classical level, there is no issue in the matter sector, as the matter field equations ultimately reduce to those identical to that obtained in the standard variational principle—there is no runaway behavior (see also [51]). In the gravitational sector, one can place restrictions on the initial data such that the system becomes dynamically equivalent to the Einstein equations with a CC. Regarding the Hamiltonian, we note that on shell, both the matter and gravitational actions vanish in our approach, and it is not too difficult to convince oneself that the total matter Hamiltonian must also vanish on shell. Though there is no issue at the classical level, the appearance of ghosts may be of particular concern for the quantization of the theory. Quantization is beyond the scope of the present article (see the discussion in [16] regarding ghosts in higher curvature theories), but we leave the reader with a couple of remarks. First, as argued in the preceding paragraph, the action presented in this article may require further modification at high energies, where the structure of the resulting action may differ significantly. Second, the vanishing of the Hamiltonian indicates that these theories suffer from the problem of time, in which the Hamiltonian operator does not yield the time evolution of the quantum state of the field, a well-known problem in canonical approaches to quantum gravity [52, *Isham:1992ms].
@@
Acknowledgments — We thank Marco de Cesare, Naresh Dadhich, and Victor Tapia for helpful comments and references. PC and JCF acknowledge support from the Taiwan National Science and Technology Council (NSTC) under project numbers 110-2112-M-002-031 and 112-2811-M-002-132, and also from the Leung Center for Cosmology and Particle Astrophysics (LeCosPA), National Taiwan University.